OpenStudy (anonymous):

Choose the correct simplification of the expression the quantity 4 times x all over y all to the second power. 4 times x to the second power all over y to the second power 8 times x to the second power all over y to the second power 16 times x to the second power all over y to the second power 16 times x to the second power all over y

OpenStudy (lgbasallote):

hint: \[\huge \left(\frac{4x}y \right)^2 \implies \frac{(4x)^2}{y^2}\] does that help?

OpenStudy (anonymous):

is that the answer?

OpenStudy (anonymous):

@dpaInc plzz help!!!!!!!!

OpenStudy (anonymous):

all you need to do now is simplify what mr. lgbassallote did... can you simplify \(\large (4x)^2 = \) ???

OpenStudy (anonymous):

how do i do that

OpenStudy (anonymous):

what does \(\large (something)^2 \) mean?

OpenStudy (anonymous):

idk

OpenStudy (anonymous):

what does \(\large a^2 \) mean ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

????

OpenStudy (anonymous):

how do i simplify 4x sq.

OpenStudy (anonymous):

\(\large a^2 \) means \(\large a \cdot a \) \(\large (something)^2 \) means \(\large (something) \cdot (something) \) so \(\large (4x)^2 \) means \(\large (4x)\cdot (4x) \)

OpenStudy (anonymous):

so the answer is D.

OpenStudy (anonymous):

no

OpenStudy (anonymous):

:( ok

OpenStudy (anonymous):

then.....

OpenStudy (anonymous):

so the answer has to be C.

OpenStudy (anonymous):

\(\huge \frac{(4x)^2}{y^2}=\frac{(4x)\cdot (4x)}{y^2}=\frac{4\cdot x \cdot4 \cdot x}{y^2} =\) ?????

OpenStudy (anonymous):

do you understand why it's C?

OpenStudy (anonymous):

yes i do thanks